OpenStudy (anonymous):

∫[e^(3t)sin(3t),dt]= ;Integral from0 to π

OpenStudy (anonymous):

You can use integration by parts take : - f' = e^3t and g = sin(3t) - f = e^3t/3 and g = 3cos(3t) use the formula now and substitute : \[\int\limits_{}^{}f'gdx = fg - \int\limits_{}^{} fg'dx\] I will let you do it, give it a try now :)

OpenStudy (anonymous):

did you get the answer almal?

OpenStudy (anonymous):

My possible answers are: 1)0; 2) (e^(3π))/2 ; 3)(1/6)(e^(3π)+1); 4)(1/6)(1-e^(3π))